An automobile with an initial velocity of 7.25 m/s accelerates at the rate of 4.0 m/s squared. What is the final velocity after 9.5 seconds?


1
Expert's answer
2021-08-16T08:33:50-0400

Given:

"v_i=7.25\;\rm m/s"

"a=4.0\:\rm m/s^2"

"t=9.5\:\rm s"


The final velocity

"v_f=v_i+at=7.25+4.0*9.5=45\:\rm m/s"